OpenStudy (parissfrance09):

help me i dont understand asap please 4^3*4^4 __________ 4^9 A) 4^16 B) 4^2 C) 1 __ 4^2 D) 1 __ 4

OpenStudy (anonymous):

order of operations, do the 4^3 and 4^4 first, then multiply them together

OpenStudy (parissfrance09):

so 3*4?

OpenStudy (anonymous):

4^3=4*4*4 4^4=4*4*4*4

OpenStudy (parissfrance09):

12 and 16

OpenStudy (anonymous):

not quite, you added them together, you want to multiply. so 4 times 4 times 4 = 4^3

OpenStudy (anonymous):

mulitplying same base to different exponent. 4^(4+3)-9. this is 4^(-2). This would be 1/16.

OpenStudy (parissfrance09):

ohhh

OpenStudy (anonymous):

AKA C, 1/4^2. Hope this helps!

OpenStudy (parissfrance09):

thanks

OpenStudy (anonymous):

\[\huge b^m\times b^n=b^{m+m}\] ile. add the exponents

OpenStudy (anonymous):

i don't see any negative exponents there, just add

OpenStudy (parissfrance09):

thank you everyone fr you guys help i think i understand now

OpenStudy (anonymous):

Good job, sorry i couldn't be more clear!

OpenStudy (parissfrance09):

you did great